OpenStudy (anonymous):

PLEASE HELP If cos(x)=sqrt(10)/10 fin csc(x)

OpenStudy (anonymous):

find sec*

OpenStudy (freckles):

\[\cos(x)=\frac{\sqrt{10}}{10}\] sec? sec is just the reciprocal of sqrt(10)/10

OpenStudy (isaiah.feynman):



OpenStudy (freckles):

if cos(x)=a/b then sec(x)=b/a if sin(x)=c/d then csc(x)=d/c if tan(x)=p/q then cot(x)=q/p by the way these if statements can also go the other away around

OpenStudy (isaiah.feynman):

Use the pythagorean theorem to find the other side, so you can get the secant of x.

OpenStudy (anonymous):

ok thanks

OpenStudy (isaiah.feynman):